# Hyperluminous Supersoft X-ray Sources in the Chandra Catalog

### v1

@@ 1 Introduction
The best-studied and best-represented subclass of HSSs is that of tidal disruption events (TDEs). TDEs occur when a star orbit happens to plunge deep enough into the potential well of a SMBH to be disrupted by tidal forces. In the standard scenario, roughly half of the resultant stellar debris remains on bound orbits and eventually falls back on the SMBHs and later is accreted after forming a bright accretion disc. The electromagnetic emission generated by this newly formed accretion disc is predicted to peak in the ultraviolet (UV) band, while the steep Wien tail of the spectrum would fall in the X-ray band. This emission is expected to rapidly rise and then slowly decay over timescales of days to months. These phenomena were first theorized (Hills, 1975; Rees, 1988; Phinney, 1989), and first observed by ROSAT (Bade et al., 1996). The properties of the first TDEs observed were exactly the ones predicted: bright and soft X-ray transients from otherwise quiescent galactic nuclei. The importance of these phenomena was immediately clear. They represent unique opportunities to study newly-formed accretion flows on otherwise quiescent and relatively “light” SMBHs ($\sim 10^{6}$ M), which are the bulk of the population of SMBHs. This is because the tidal radius for a Sun-like star is smaller than the Schwarzschild radius for a SMBH with a mass $\gtrsim 10^{8}$ M, hence a star would not get disrupted, but rather directly swallowed by any more massive SMBH: one such encounter would not generate a TDE.
@@
After the first TDEs were discovered in the X-ray band, these phenomena were found to be bright across the full electromagnetic spectrum, and today we know of about a hundred of TDEs observed in every wavelength, from radio (Cendes et al., 2024), to infrared (Masterson et al., 2024), optical/UV (Gezari et al., 2012; van Velzen et al., 2021; Hammerstein et al., 2023), X-ray (Saxton et al., 2020; Sazonov et al., 2021), and Gamma-ray (Burrows et al., 2011). The simple picture described above struggles to explain all of the data accumulated on these phenomena in the last three decades. Ongoing debates revolve around where, when, and how the multi-wavelength properties of TDEs originate and are correlated (Dai et al., 2018; Bonnerot et al., 2021, e.g.). To further complicate this scenario, new classes of bright and supersoft transients from SMBHs have emerged in recent years. The discovery of quasi-periodic eruptions (QPEs), spectacular X-ray flashes from galactic nuclei recurring over hourly time scales(Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021), repeated/partial TDEs (Payne et al., 2021; Wevers et al., 2023; Malyali et al., 2023; Lin et al., 2024; Guolo et al., 2024), long-lived TDEs (Lin et al., 2017b; He et al., 2021a; Goodwin et al., 2022) and active galactic nuclei (AGN) in a purely thermal state (He et al., 2021b; Lin et al., 2017a), calls for further theoretical and observational efforts in order to explain the nature and characteristic of these phenomena. The properties of this plethora of sources are similar enough to those of “classical” TDEs that a comprehensive model of the latter should also be able to somewhat account for the former. Although all of these classes of sources exhibit different temporal evolutions, their X-ray properties, in terms of spectrum and overall luminosity, are strikingly similar; hence we collect them all and define the HSS category.
@@ 2.1 Sample selection
Figure 1: X-ray luminosity as a function of the HR in the soft-to-medium band for the source selected via their X-ray properties. In black are the sources we then excluded based on their optical classification, in blue are the HSSs known in the literature, and in red are the sources classified as non-active galaxies whose X-ray spectrum we finally analyzed.
@@
For all detections, we retrieved their fluxes, hardness ratios (HRs), significances in different energy bands, and errors in each of these quantities. We selected our sources based on their significance (S) in the soft (0.5–1.2 keV) band, HR in the soft to medium (1.2–2 keV) band, and luminosity in the broad (0.5–7 keV) band. We note that, given that the emission from HSSs is extremely soft, the natural choice would be to employ also the ultra-soft (0.2–0.5 keV) band. However, the progressive and dramatic decrease in effective area of Chandra’s detectors, especially at ultra-soft energies (Plucinsky et al., 2018, 2022), implies that the ultra-soft band is not reliable for source selection. We compute source luminosities from the X-ray fluxes using redshift measurements. We crossmatched the Chandra catalog with all the redshift catalogs available in SIMBAD (Wenger et al., 2000). We performed the crossmatch using the tool TOPCAT (Taylor, 2005), with a matching radius corresponding to the error on the source position. Knowing the redshift, the luminosity values are computed by assuming a standard flat $\Lambda$CDM cosmology with $\Omega_{\textup{M}}=0.3$ and $H_{0}=70$ km s-1 Mpc-1.
@@
To build our sample we discarded all sources with a broad-band luminosity $L_{\textup{X}}<10^{41}$ erg/s, HR $>-0.55$ in the soft-to-medium band and S $>7$ in the soft band. The choice of HR corresponds to a photon index $\Gamma\gtrsim 3$ if the X-ray spectra of the sources were modeled with a power-law profile. These three criteria ensure that we are excluding sources powered by stellar processes (which rarely exceed the chosen luminosity threshold), the bulk of the AGN population (which usually exhibits much flatter spectra), and that the source is significant enough to perform a reliable spectral and timing analysis. This resulted in 222 detections from 125 individual sources. Figure 1 shows the X-ray luminosity as a function of the HR for all selected detections.
@@
These 222 detections represent about four times as many HSS candidates as were present in the XMM catalog from Sa23, even though the initial catalogs are similarly-sized. This is because we are not imposing any conditions on the ultra-soft (0.2–0.5 keV) band, and hence we expect more spurious sources to be present in this sample. To further clean our selection we thus exploit the optical classification of the sources, which we retrieved along with the redshift information from SIMBAD. Straightaway we identified and excluded 24 spurious sources: ten sources associated with stars, one located behind a supernova remnant, four associated with a foreground galaxy, seven whose soft X-ray photons are associated with the emission from galaxy clusters, and two whose emission is extended and attributable to a pair of interacting galaxies. Similarly to what we found in Sa23, at this stage, the bulk of our sample is composed of sources optically classified as AGN, 75 as broad-line AGN, and 15 Seyfert 2 galaxies (mostly classified by Véron-Cetty & Véron 2010 and Pâris et al. 2014). The X-ray emission of these sources passes our supersoft criterion as it is characterized by a strong soft excess dominating the energy range below 2 keV. However, they also exhibit hard X-ray emission, which can be well-reproduced by a power-law model (seen through different ranges of absorption material). Overall, the X-ray emission of these sources is compatible with their optical classification and we will not address them further.
@@ 2.2 X-ray spectral analysis
All Chandra datasets were reprocessed and reduced with the Chandra Interactive Analysis of Observations software package (CIAO, v.4.12; Fruscione et al. 2006) and the CALDB 4.9.0 release of the calibration files. For each source, counts were extracted using the CIAO tool srcflux, from regions encompassing a point-spread function (PSF) fraction of about 95%. Local backgrounds were estimated from source-free annular regions, centered on the source position. Spectra, spectral redistribution matrices, and ancillary response files were generated using the CIAO script specextract. The spectra were then fed into the spectral fitting package XSPEC (Arnaud, 1996) version 12.12.1.
@@ 3 A new “old” TDE
The X-ray spectrum of J1035+57, reported in Fig. 3, can be well reproduced by a simple black-body model (redshifted at the source distance, zbbody in XSPEC). Given the high signal-to-noise ratio of the detection, we rebinned the spectrum to have at least 20 counts in each bin and adopted the standard $\chi^{2}$ statistic. Considering the supersoft nature of the source’s spectrum and the fact that the dataset was acquired in 2002 when Chandra effective area in the soft X-ray band was not degraded by molecular contamination, we also include the counts collected in the 0.3–0.5 keV band in our analysis. To the black-body model, we added a layer of neutral absorber (TBabs) with $N_{\textup{H}}$ fixed to the Galactic value $5.46\times 10^{19}$ cm-2 (HI4PI Collaboration et al., 2016), using abundances from Wilms et al. (2000), with the photoelectric absorption cross-sections from Verner et al. (1996). We obtained a satisfactory fit ($\chi^{2}/\nu=36.3/29\approx 1.24$, where $\chi^{2}$ is the value of the statistic and $\nu$ the degrees of freedom) for a temperature of $85\pm 3$ eV and an unabsorbed luminosity of $1.28\pm 0.07\times 10^{42}$ erg/s in the 0.5–2 keV band. The spectrum residuals offer hints of excess counts in the 1–2 keV band, but adding a power law to reproduce this excess does not significantly improve the fit’s quality. Likewise, the quality of the fit is not improved by adding a free-to-vary layer of neutral absorption. The 3-$\sigma$ upper limit on the intrinsic absorption at the source location (or within the host galaxy) is $N_{\textup{H,intrinsic}}<3\times 10^{20}$ cm-2. We also attempted fitting the spectrum with a power-law and an apec model, and in both cases, we obtained unacceptable fits ($\chi^{2}/\nu=57.5/29\approx 1.98$ and $\chi^{2}/\nu=127.9/29\approx 4.41$ respectively).
Figure 3: X-ray spectrum (upper panels) and residuals (lower panels) of J1035+57. The solid lines show the best-fitting models described in the text.
@@
The host galaxy properties, X-ray source location, long-term variability, spectral shape, and luminosity, all suggest that J1035+57 emission is due to a TDE. The non-detection dating one year before the detection places the date of the TDE in the second half of 2001 or the first half of 2002, making this event one of the first TDEs ever observed by Chandra.
@@ 4 Machine-learning approach
In the previous sections, we manually applied the algorithm developed in Sa23 and successfully identified a previously-overlooked TDE in the Chandra catalog. In this section, we explore the applicability and efficiency of ML techniques for this task.
In machine learning, representation learning(Bengio et al., 2013), refers to the process of training a deep neural network to perform a given task (e.g. spectral or light curve reconstruction), by learning a low-dimensionality representation of the input data. This low-dimensional learned representation can then be used for other downstream tasks, such as classification or regression on relevant parameters. It can also be used for anomaly detection in large datasets characterized by high dimensionality and represents a more flexible and scalable approach with respect to classical methods.
In the X-ray band, the application of classification algorithms based on machine learning techniques has only been recently explored (Yang et al., 2022; Pérez-Díaz et al., 2024), with most efforts focusing on the identification of variability patterns, and in particular of fast X-ray transients (Dillmann et al., 2024, 2025). Representation learning algorithms build compressed, low-dimensional representations of high-dimensionality objects. In these learned embedding spaces, anomalies are typically isolated from the bulk of ”normal” objects, and HSSs, the targets of our search, owing to their spectral properties, are expected to be anomalies in this representation space.
@@
This approach, which led to the discovery of at least one new fast extragalactic transient (Dillmann et al., 2024), was trained using 95,473 Chandra filtered event files corresponding to 58,932 sources. This is a smaller dataset than the one we analyzed in the previous section, but notably, it included three HSSs: SDSS J150052.07+015453.8, WINGS J134849.88+263557.5, and two observations of ASASSN-14li. We investigate if the learned representations of these known TDEs are isolated in the representation space with respect to the rest of the X-ray sources. If so, this would indicate the algorithm’s ability to identify new HSS candidates.
@@
Figure 4 shows a visualization of the autencoder latent space after training. In addition to the autoencoder, the high-dimensional input data (a vector representing the two-dimensional distribution of photon arrival times and energies), is further reduced to the shown 2D representation using a t-distributed stochastic neighbor embedding (t-SNE, van der Maaten & Hinton 2008). The embeddings are color-coded by spectral hardness, and we have highlighted the embedding locations of the known HSSs with the black circles. Three out of the four HSSs’ observations cluster in the same “island” of the two-dimensional map. In this pilot study, we focus on this region of the embedding, which might harbor additional HSS candidates.
@@ Acknowledgments
A.S. acknowledges support from a Scholarly Studies Award. PK was supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51534.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This research has made use of data obtained from the Chandra Data Archive and the Chandra Source Catalog, both provided by the Chandra X-ray Center (CXC). This paper employs a list of Chandra data sets, obtained by the Chandra X-ray Observatory, available at https://doi.org/10.25574/cdc.334 (catalog DOI: 10.25574/cdc.334).

### v2

@@ 1 Introduction
The best-studied and best-represented subclass of HSSs is that of tidal disruption events (TDEs). TDEs occur when a star orbit happens to plunge deep enough into the potential well of a SMBH to be disrupted by tidal forces. In the standard scenario, roughly half of the resultant stellar debris remains on bound orbits and eventually falls back on the SMBHs and later is accreted after forming a bright accretion disc. The electromagnetic emission generated by this newly formed accretion disc is predicted to peak in the ultraviolet (UV) band, while the steep Wien tail of the spectrum would fall in the X-ray band. This emission is expected to rapidly rise and then slowly decay over timescales of days to months. These phenomena were first theorized (Hills, 1975; Rees, 1988; Phinney, 1989), and first observed by ROSAT (Bade et al., 1996). The properties of the first TDEs observed were exactly the ones predicted: bright and soft X-ray transients from otherwise quiescent galactic nuclei. The importance of these phenomena was immediately clear. They represent unique opportunities to study newly-formed accretion flows on otherwise quiescent and relatively “light” SMBHs ($\sim 10^{6}$ M). This is because the tidal radius for a Sun-like star is smaller than the Schwarzschild radius for a SMBH with a mass $\gtrsim 10^{8}$ M, hence a star would not get disrupted, but rather directly swallowed by any more massive SMBH: one such encounter would not generate a TDE.
@@
After the first TDEs were discovered in the X-ray band, these phenomena were found to be bright across the full electromagnetic spectrum, and today we know of about a hundred of TDEs observed in every wavelength, from radio (Cendes et al., 2024), to infrared (Masterson et al., 2024), optical/UV (Gezari et al., 2012; van Velzen et al., 2021; Hammerstein et al., 2023), X-ray (Saxton et al., 2020; Sazonov et al., 2021), and Gamma-ray (Burrows et al., 2011). The simple picture described above struggles to explain all of the data accumulated on these phenomena in the last three decades. Ongoing debates revolve around where, when, and how the multi-wavelength properties of TDEs originate and are correlated (Dai et al., 2018; Bonnerot et al., 2021, e.g.). To further complicate this scenario, new classes of bright and supersoft transients from SMBHs have emerged in recent years. The discovery of quasi-periodic eruptions (QPEs), spectacular X-ray flashes from galactic nuclei recurring over hourly time scales(Miniutti et al., 2019; Giustini et al., 2020; Arcodia et al., 2021; Quintin et al., 2023; Nicholl et al., 2024), repeated/partial TDEs (Payne et al., 2021; Wevers et al., 2023; Malyali et al., 2023; Lin et al., 2024; Guolo et al., 2024), long-lived TDEs (Lin et al., 2017b; He et al., 2021a; Goodwin et al., 2022) and active galactic nuclei (AGN) in a purely thermal state (He et al., 2021b; Lin et al., 2017a), calls for further theoretical and observational efforts in order to explain the nature and characteristic of these phenomena. The properties of this plethora of sources are similar enough to those of “classical” TDEs that a comprehensive model of the latter should also be able to somewhat account for the former. Although all of these classes of sources exhibit different temporal evolutions, their X-ray properties, in terms of spectrum and overall luminosity, are strikingly similar; hence we collect them all and define the HSS category.
@@ 2.1 Sample selection
Figure 1: X-ray luminosity as a function of the HR in the soft-to-medium band for the source selected via their X-ray properties. In black are the sources we then excluded based on their optical classification, in blue are the HSSs known in the literature, and in red are the sources classified as non-active galaxies whose X-ray spectrum we finally analyzed. The error bars correspond to $1\sigma$ uncertainties.
@@
For all detections, we retrieved their fluxes, hardness ratios (HRs), significances (Ss) in different energy bands, and errors in each of these quantities. We adopted aperture flux to be as independent as possible from particular choices of spectral models, and flux significances, which is a simple estimate of the ratio of the flux measurement to its average error. We selected our sources based on their S in the soft (0.5–1.2 keV) band, HR in the soft to medium (1.2–2 keV) band, and luminosity in the broad (0.5–7 keV) band. We compute the absorbed luminosities from the X-ray fluxes using redshift measurements. We crossmatched the Chandra catalog with all the redshift catalogs available in SIMBAD (Wenger et al., 2000). We performed the crossmatch using the tool TOPCAT (Taylor, 2005), with a matching radius corresponding to the error on the source position (provide by the Chandra catalog at 90% confidence level), and considering the best optical match for each X-ray source, as multiple X-ray sources might be hosted in the same galaxy. Knowing the redshift, the luminosity values are computed by assuming a standard flat $\Lambda$CDM cosmology with $\Omega_{\textup{M}}=0.3$ and $H_{0}=70$ km s-1 Mpc-1. We note that, given that the emission from HSSs is extremely soft, the natural choice would be to employ also the ultra-soft (0.2–0.5 keV) band. However, the progressive and dramatic decrease in effective area of Chandra’s detectors, especially at ultra-soft energies (Plucinsky et al., 2018, 2022), implies that the ultra-soft band is not reliable for source selection.
@@
To build our sample we discarded all sources with a broad-band luminosity $L_{\textup{X}}<10^{41}$ erg/s, HR $>-0.55$ in the soft-to-medium band and S $>7$ in the soft band. The choice of HR corresponds to a photon index $\Gamma\gtrsim 3$ if the X-ray spectra of the sources were modeled with a power-law profile. These three criteria ensure that we are excluding sources powered by stellar processes (which rarely exceed the chosen luminosity threshold), the bulk of the AGN population (which usually exhibits much flatter spectra), and that the source statistics is high enough to perform a reliable spectral and timing analysis. This resulted in 222 detections from 125 individual sources. Figure 1 shows the X-ray luminosity as a function of the HR for all selected detections.
@@
These 222 detections represent about four times as many HSS candidates as were present in the XMM catalog from Sa23, even though the initial catalogs are similarly-sized. This is because we are not imposing any conditions on the ultra-soft (0.2–0.5 keV) band, and hence we expect more spurious sources to be present in this sample. To further clean our selection we thus exploit the optical classification of the sources, which we retrieved along with the redshift information from SIMBAD. Straightaway we identified and excluded 24 spurious sources (86 detections): ten sources associated with stars, one located behind a supernova remnant, four associated with a foreground galaxy, seven whose soft X-ray photons are associated with the emission from galaxy clusters, and two whose emission is extended and attributable to a pair of interacting galaxies. Similarly to what we found in Sa23, at this stage, the bulk of our sample is composed of sources optically classified as AGN, 75 (96 detections) as broad-line AGN, and 15 (24 detections) Seyfert 2 galaxies (mostly classified by Véron-Cetty & Véron 2010 and Pâris et al. 2014). The X-ray emission of these sources passes our supersoft criterion as it is characterized by a strong soft excess dominating the energy range below 2 keV. However, they also exhibit hard X-ray emission, which can be well-reproduced by a power-law model (seen through different ranges of absorption material). Overall, the X-ray emission of these sources is compatible with their optical classification and we will not address them further.
@@ 2.2 X-ray spectral analysis
All Chandra datasets were reprocessed and reduced with the Chandra Interactive Analysis of Observations software package (CIAO, v.4.12; Fruscione et al. 2006) and the CALDB 4.9.0 release of the calibration files. For each source, spectra, spectral redistribution matrices, and ancillary response files were generated using the CIAO script specextract. Spectra were extracted from regions encompassing a point-spread function (PSF) fraction of about 95%. Local backgrounds were estimated from source-free annular regions, centered on the source position. These annular regions have inner radii of $\approx 10$ arcsec and outer radii of $\approx 30$ arcsec. If these regions are too close to the detector borders, a circular region of $\approx 20$ arcsec radius was adopted instead.
@@ 3 A new “old” TDE
The X-ray spectrum of J1035+57, reported in Fig. 3, can be well reproduced by a simple black-body model (redshifted at the source distance, zbbody in XSPEC). Given the high signal-to-noise ratio of the detection, we rebinned the spectrum to have at least 20 counts in each bin and adopted the standard $\chi^{2}$ statistic. Considering the supersoft nature of the source’s spectrum and the fact that the dataset was acquired in 2002 when Chandra effective area in the soft X-ray band was not degraded by molecular contamination, we also include the counts collected in the 0.3–0.5 keV band in our analysis. To the black-body model, we added a layer of neutral absorber (TBabs) with $N_{\textup{H}}$ fixed to the Galactic value $5.46\times 10^{19}$ cm-2 (HI4PI Collaboration et al., 2016), using abundances from Wilms et al. (2000), with the photoelectric absorption cross-sections from Verner et al. (1996). We obtained a satisfactory fit ($\chi^{2}/\nu=36.3/29\approx 1.24$, where $\chi^{2}$ is the value of the statistic and $\nu$ the degrees of freedom) for a temperature of $85\pm 3$ eV and an unabsorbed luminosity of $1.28\pm 0.07\times 10^{42}$ erg/s in the 0.5–2 keV band. The spectrum residuals offer hints of excess counts in the 1–2 keV band, but adding a power law to reproduce this excess does not significantly improve the fit’s quality. Likewise, the quality of the fit is not improved by adding a free-to-vary layer of neutral absorption. The 3-$\sigma$ upper limit on the intrinsic absorption at the source location (or within the host galaxy) is $N_{\textup{H,intrinsic}}<3\times 10^{20}$ cm-2. We also attempted fitting the spectrum with alternative models, such as a power-law and an apec model, and in both cases, we obtained unacceptable fits ($\chi^{2}/\nu=57.5/29\approx 1.98$ and $\chi^{2}/\nu=127.9/29\approx 4.41$ respectively).
Figure 3: X-ray spectrum (upper panels) and residuals (lower panels) of J1035+57. The solid lines show the best-fitting models described in the text. Only the 0.3–2 keV energy range is shown as above 2 keV the spectrum is background dominated.
@@
The host galaxy properties, X-ray source location, long-term variability, spectral shape, and luminosity, all suggest that J1035+57 emission is due to a TDE. The non-detection dating one year before the detection places the date of the TDE in the second half of 2001 or the first half of 2002.
@@ 4 Machine-learning approach
In the previous sections, we manually applied the algorithm developed in Sa23 and successfully identified a previously-overlooked TDE in the Chandra catalog. In this section, we explore the applicability and efficiency of a ML technique developed in Dillmann et al. (2025) for this task.
In ML, representation learning (Bengio et al., 2013), refers to the process of training a deep neural network to learn a low-dimensional informative representation of the input data that can be used for various downstream tasks, such as classification or regression on relevant parameters. It can also be used for similarity searches and anomaly detection in large datasets characterized by high dimensionality and represents a more flexible and scalable approach with respect to classical methods.
In the X-ray band, the application of classification algorithms based on ML techniques has only been recently explored (Yang et al., 2022; Pérez-Díaz et al., 2024). Most efforts have focused on the identification of variability patterns, and in particular of fast X-ray transients (Dillmann et al., 2025, 2025). Representation learning algorithms build compressed, low-dimensional representations of high-dimensionality objects. In these learned embedding spaces, anomalies are typically isolated from the bulk of “normal” objects, and HSSs, the targets of our search, owing to their spectral properties, are expected to be anomalies in this representation space.
@@
This approach, which led to the discovery of the new fast extragalactic transient XRT 200515 (Dillmann et al., 2025), was trained using 95,473 Chandra filtered event files corresponding to 58,932 sources. This is a smaller dataset than the one we analyzed in the previous section, but notably, it included three HSSs: SDSS J150052.07+015453.8, WINGS J134849.88+263557.5, and two observations of ASASSN-14li. We investigate if the learned representations of these known TDEs are isolated in the representation space with respect to the rest of the X-ray sources. If so, this would indicate the algorithm’s ability to identify new HSS candidates.
@@
Figure 4 shows a visualization of the autoencoder latent space after training. In addition to the autoencoder, the high-dimensional input data (a vector representing the two-dimensional distribution of photon arrival times and energies), is further reduced to the shown 2D representation using a t-distributed stochastic neighbor embedding (t-SNE, van der Maaten & Hinton 2008). The embeddings are color-coded by spectral hardness, and we have highlighted the embedding locations of the known HSSs with the black circles. Three out of the four HSSs’ observations, the two of ASASSN-14li and the single of WINGS J134849.88+263557.5, cluster in the same “island” of the two-dimensional map. In this pilot study, we focus on using this representation learning approach for similarity searches around the known HSSs. The embedding region in which they live is well-isolated and might harbor additional HSS candidates.
@@ Acknowledgments
We thank the anonymous referee for insightful comments that improved the paper. A.S. acknowledges support from a Scholarly Studies Award. PK was supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51534.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555. This research has made use of data obtained from the Chandra Data Archive and the Chandra Source Catalog, both provided by the Chandra X-ray Center (CXC). This paper employs a list of Chandra data sets, obtained by the Chandra X-ray Observatory, available at https://doi.org/10.25574/cdc.334 (catalog DOI: 10.25574/cdc.334).